OpenStudy (anonymous):

Coral created a painting that was 4 1/3 ft long and 5 1/5 ft wide. What was the area of Coral's painting? Express your answer in simplest form

OpenStudy (anonymous):

\[\large \sf A=l \times w\] \[\large \sf A=(4+\frac{1}{3})(5+\frac{1}{5})\] Just distribute aka foil method.

OpenStudy (anonymous):

wut?

RhondaSommer (rhondasommer):

i can explain it

RhondaSommer (rhondasommer):

how do you get the are of something?

RhondaSommer (rhondasommer):

area*

OpenStudy (anonymous):

You went from 13/3 to 13/5

RhondaSommer (rhondasommer):

how do you get area @LegendarySadist

OpenStudy (anonymous):

\[\large \sf A=l \times w\]

OpenStudy (anonymous):

length, not height

RhondaSommer (rhondasommer):

sorry, simple mistake :)

RhondaSommer (rhondasommer):

multiply width times length

OpenStudy (anonymous):

For this kind of multiplication I prefer the foil method. \[\large \sf (4+\frac{1}{3})(5+\frac{1}{5})=(4\times 5)+(4 \times \frac{1}{5})+(\frac{1}{3} \times 5)+(\frac{1}{3} \times \frac{1}{5})\]

RhondaSommer (rhondasommer):

that could work too!

RhondaSommer (rhondasommer):

something more simple that you could do is \[4*5 . and . 1/3*1/5 add . them . together :)\]

OpenStudy (anonymous):

That wouldn't work.

RhondaSommer (rhondasommer):

your right sorry...dont know what i was typing lol

RhondaSommer (rhondasommer):

what you like to do would work best...I was thinking of a different type of math.

OpenStudy (anonymous):

i dont get nth